# Bayesian Functional Graphical Models with Change-Point Detection

### v1

@@ Abstract
Functional data analysis, which models data as realizations of random functions over a continuum, has emerged as a useful tool for time series data. Often, the goal is to infer the dynamic connections (or time-varying conditional dependencies) among multiple functions or time series. For this task, we propose a dynamic and Bayesian functional graphical model. Our modeling approach prioritizes the careful definition of an appropriate graph to identify both time-invariant and time-varying connectivity patterns. We introduce a novel block-structured sparsity prior paired with a finite basis expansion, which together yield effective shrinkage and graph selection with efficient computations via a Gibbs sampling algorithm. Crucially, the model includes (one or more) graph changepoints, which are learned jointly with all model parameters and incorporate graph dynamics. Simulation studies demonstrate excellent graph selection capabilities, with significant improvements over competing methods. We apply the proposed approach to study of dynamic connectivity patterns of sea surface temperatures in the Pacific Ocean and discovers meaningful edges.
@@ 1 Introduction
The main tool for inferring conditional dependencies among random functions is functional graphical models (FGMs), where each node represents an entire function instead of a scalar variable. Like for traditional (non-functional) graphical models, there are primarily two approaches for graph estimation with FGMs. The first approach is neighborhood selection, which regresses each variable on all other variables and identifies conditional dependencies based on nonzero regression coefficients. Zapata et al., (2022) and Zhao et al., (2021) adapt this approach for functional data using function-on-function regressions, i.e., linear regression models with a functional response and functional covariates. Related, Wu et al., (2022) estimate multiple graphs simultaneously and add a Frobenius norm penalty to enforce graph similarity. Nonparametric variations that relax linearity and distributional assumptions (see below) are also available (Li and Solea,, 2018; Solea and Dette,, 2022; Solea and Li,, 2022; Lee et al.,, 2022) and may incorporate external covariates (Lee et al.,, 2021). However, these methods are exclusively non-Bayesian, do not provide uncertainty quantification, and typically require the selection of multiple tuning parameters.
@@
In general, these methods only estimate a single, time-invariant graph for the multivariate functional data. However, when the data are functions of time, it is often of interest to estimate a dynamic graph, so that the conditional dependencies may change through time. For instance, connectivity patterns among SSTs may differ depending on the phase of the seasonal cycle, motivating dynamic graph estimation for multivariate functional data. This generalization requires careful definition of an appropriate graph. Qiao et al., (2020) propose dynamic graph estimation by (i) smoothing the functional data and (ii) separately estimating sparse graphs at each time $t$. Nodes in this dynamic graph at time $t$ represent functions evaluated only at time $t$. However, the graph at time $t$ is influenced by the functional data observed on the whole time domain as a result of the smoothing in (i), which confuses the interpretation of the dynamic graph and may introduce sensitivity to data far from $t$. Further, this approach requires selection of multiple tuning parameters, including a sparsity parameter for every time $t$, and does not provide uncertainty quantification. Alternatively, Zhang et al., (2021) introduce a Bayesian model that defines an edge when two functions are conditionally dependent at time $t$ given all other functions at all other times. Clearly, this graph definition deviates from traditional and functional graphical models: the conditional dependence here does not correspond exactly to the nodes in the graph. Further, Zhang et al., (2021) require local, basis coefficient-specific shrinkage priors, which can induce unnecessary variability unless additional structure is imposed (Gao and Kowal,, 2022).
@@ 2.1 Bayesian functional graphical models
The goal of functional graphical models (FGMs) is to learn and express the conditional dependencies between random functions. Unlike traditional graphical models, where the nodes usually represent scalar random variables, the FGM nodes represent random functions on their full domain. Let $\bm{Y}^{i}=\{Y^{i}_{j}(t)\}_{j=1}^{p}$ denote iid replicates for $i=1,\ldots,n$ of $p$ random functions $\{y_{j}\}_{j=1}^{p}$ on the functional domain $t\in\mathcal{T}\subseteq\mathbb{R}$. Since these functions may be observed subject to measurement error, we similarly define $\{X_{j}^{i}(t)\}_{j=1}^{p}$ to be iid realizations of the error-free functions $\{x_{j}\}_{j=1}^{p}$. The conditional dependencies among the $p$ random functions are represented by an undirected graph $\mathcal{G}(V,E)$ with vertices $\mathcal{V}=\{1,\ldots,p\}$ and undirected edges $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$. An edge $(j_{1},j_{2})$ exist if and only if the corresponding two functions are not conditionally independent. Specifically, we define conditional independence for functional data as in Zhu et al., (2016): letting $A,B,C\subset\{1,\ldots,p\}$ we say that $\bm{x}_{A}$ and $\bm{x}_{B}$ are conditionally independent given $\bm{x}_{C}$ if and only if the conditional distribution $[\bm{x}_{A}\mid\bm{x}_{B},\bm{x}_{C}]$ equals the conditional distribution $[\bm{x}_{A}\mid\bm{x}_{C}]$.
@@
The fundamental idea of Gaussian graphical models is that conditional independence corresponds to sparsity in the precision matrix. Let $\bm{\Omega}_{j_{1},j_{2}}=\{\omega_{j_{1}k_{1},j_{2}k_{2}}\}_{k_{1},k_{2}=1}^{K}$ denote the $(j_{1},j_{2})$th block of $\bm{\Omega}$ for $j_{1},j_{2}=1,\ldots,p$, so each block is $K\times K$. Under the Gaussianity assumption (3), it is easy to show that a zero block $\bm{\Omega}_{j_{1},j_{2}}=\bm{0}_{K\times K}$ is equivalent to the conditional independence between $\bm{c}_{j_{1}}$ and $\bm{c}_{j_{2}}$ in (4). Therefore, we deduce that, based on (4), a zero block $\bm{\Omega}_{j_{1},j_{2}}=\bm{0}_{K\times K}$ is equivalent to the absence of an edge between nodes $j_{1}$ and $j_{2}$ in the functional space $\mathcal{G}(V,E)$. In other words, learning the edge set $E$ of the graph $\mathcal{G}(V,E)$ is equivalent to learning the nonzero entries of $\bm{\Omega}$. In order to recover edges of the graph $\mathcal{G}$ in the functional space, we must learn the block-sparsity structure of $\bm{\Omega}$.
@@
Crucially, this prior incorporates the block-wise structure: if $\pi_{j_{1}j_{2}}$ is close to zero, the edges in the $(j_{1},j_{2})$th block of $\bm{G}$ are very likely to be all zeros. As a result, the edge $(j_{1},j_{2})$ in the functional space graph $\mathcal{G}$ will not be included. Section 2.2.3 provides supplementary analysis of the prior based on Monte Carlo samples. The hyperparameters $\alpha$ and $\beta$ in the Beta distribution control the level of sparsity in the graph $\mathcal{G}$. When the mean is higher, the resulting graph in the functional space is denser. In the simulation study and the real data analysis, we assume that the on-diagonal blocks are always fully connected and fix $\pi_{0}=0$ for simplicity; results are not sensitive to this choice. Lastly, we specify an inverse-Gamma prior for $\sigma_{\epsilon}^{2}$ with shape $\alpha_{\sigma}$ and rate $\beta_{\sigma}$.
@@ 2.2.3 Choice of hyperparameters
Bayesian graphical models commonly set the prior marginal edge inclusion probability to $2/(p-1)$, which implies that the prior expected number of edges to $p$. Accordingly, we set the prior belief of edge inclusion probabilities of graph $\bm{G}^{(s)}$ in the coefficient space to $2/(p-1)$, corresponding to an expected number of edges equal to $pK^{2}$. As in Wang, (2015) and Peterson et al., (2020), the prior edge inclusion probability is not analytically available because of intractable normalizing constants. A complete form of the full prior with the normalizing constants is included in Appendix A. In order to investigate whether the normalizing constants dominate the prior and drive the prior to the extreme, we evaluate the prior edge inclusion probability based on Monte Carlo sampling from the prior.
Figure 1: Prior edge inclusion probability (y-axis) versus the prior mean of $\pi_{j_{1},j_{2}}$ (x-axis). The colored lines refer to different values of $p$. Smaller $p$ corresponds to a lower line. The solid black line is y = x. The point (a,b) on the line of $p$ indicates $\alpha/(\alpha+\beta)=a/(p-1)$ and the posterior edge inclusion probability is $b/(p-1).$
First, we evaluate the influence of $\alpha$ and $\beta$ on the marginal prior edge inclusion probability. We fix $\lambda=1$, $\nu_{0}=0.02$ and $h=50$ ($v_{1}=hv_{0}$) as commonly recommended (Wang,, 2015; Peterson et al.,, 2020; Osborne et al.,, 2022). We also fix $\alpha=2$ and $K=5$. The value of $\beta$ varies so that the prior mean of the beta distribution kernel for each $\pi_{j_{1},j_{2}}$ in (7) takes values in $\{1/(p-1),2/(p-1),3/(p-1),4/(p-1),5/(p-1)\}$. The study is carried out on $p=(5,10,15,20)$. Figure 1 shows the prior edge inclusion probabilities estimated via Monte Carlo sampling. We conclude that: 1) the prior edge inclusion probability is not dominated by the normalizing constant; 2) similar to Wang, (2015), the prior edge inclusion probability is lower than the prior mean of $\pi_{j_{1},j_{2}}$; and 3) the influence of normalizing constants on the prior edge inclusion probability is greater when $p$ is smaller, but the differences among different values of $p$ are negligible. Figure 1 can be used as a guide to choose parameters that reflect the preferred edge inclusion probability. For example, the prior edge inclusion probability is roughly $2/(p-1)$ when the prior mean of $\pi_{j_{1},j_{2}}$ is $3/(p-1)$. Therefore, when $p=15$, $\alpha=2$ and $\beta=7$ is a reasonable choice.
@@ 3.3 Parameter settings
We use $K=5$ B-splines of order 3 to fit the model, which provides accurate estimation of the underlying true smooth curves. Note that this basis differs from the polynomial basis used to generate the data. For the hyperparameters in the joint prior (5), we follow the guidelines in Section 2.2.3 and set $\lambda=1$, $v_{0}=0.02$, $h=50$ ($v_{1}=hv_{0}$) $\alpha=2$ and $\beta=7$. To initialize the MCMC, the precision matrix and graphs are all set to the identity matrix and the block-wise edge inclusion probabilities are set to 0.5. Our simulation study shows that the model is robust to the initialization of the basis coefficients and the changepoint $\tau$. Results reported below were obtained by running MCMC chains for 5000 total iterations with a burn-in of 3000 iterations. The models took on average about 4 minutes on a Mac computer with a 16 GB Apple M1 Pro chip. MCMC convergence was checked by the Geweke diagnostic test (Geweke et al.,, 1991) on the key parameters $\{\bm{c}^{i(s)},\sigma_{\epsilon}^{(s)},\tau\}$.
@@ 3.5 Learning the block-sparsity
As suggested by our simulation results, our proposed model, together with the choice of hyper-parameters, leads to sparse estimation of the $p$-by-$p$ graphs in the functional space. Our prior on $\bm{G}$ does not specify specific edges in $\bm{G}$ to be zeros and ones, but rather gives a common marginal probability of inclusion to all edges. As data are added, the block-wise structure, if there is any, is reinforced by the block-specific inclusion probabilities $\pi_{j_{1}j_{2}}$. These parameters induce group shrinkage and enable information sharing among edges within a block. If the data indicate that most of the edges in a block are zero, then we expect $\pi_{j_{1}j_{2}}$ to be small, which will further discourage the inclusion of other edges in that block.
@@
$p(g_{j_{1}k_{1},j_{2}k_{2}}=1\mid\bm{\pi},\bm{\Omega})=\dfrac{N(\omega_{j_{1}k_{1},j_{2}k_{2}}\mid 0,v_{1}^{2})\pi_{j_{1}k_{1},j_{2}k_{2}}}{N(\omega_{j_{1}k_{1},j_{2}k_{2}}\mid 0,v_{1}^{2})\pi_{j_{1}j_{2}}+N(\omega_{j_{1}k_{1},j_{2}k_{2}}\mid 0,v_{0}^{2})(1-\pi_{j_{1}j_{2}})},$
which is a combination of $\pi_{j_{1},j_{2}}$ and the likelihood. In our simulation study, the posterior mean of $\pi_{j_{1},j_{2}}$ was about 0.19 for the zero blocks and 0.08 for the nonzero blocks. Edges with a high likelihood in the slab component have a high posterior marginal inclusion probability, which results in an estimated edge in the functional space. To illustrate this point, we calculated the Frobenuius norm of estimated precision matrix ^$\bm{\hat{}}{\bm{\Omega}}$ in each block, and compared the distribution of these Frobenius norms between zero and nonzero blocks (see Appendix B). There is a clear separation between the two groups. The density from the zero blocks is highly concentrated around zero, especially compared to the density from the nonzero blocks.

### v2

@@ Abstract
Functional data analysis, which models data as realizations of random functions over a continuum, has emerged as a useful tool for time series data. Often, the goal is to infer the dynamic connections (or time-varying conditional dependencies) among multiple functions or time series. For this task, a dynamic and Bayesian functional graphical model is introduced. The proposed modeling approach prioritizes the careful definition of an appropriate graph to identify both time-invariant and time-varying connectivity patterns. A novel block-structured sparsity prior is paired with a finite basis expansion, which together yield effective shrinkage and graph selection with efficient computations via a Gibbs sampling algorithm. Crucially, the model includes (one or more) graph changepoints, which are learned jointly with all model parameters and incorporate graph dynamics. Simulation studies demonstrate excellent graph selection capabilities, with significant improvements over competing methods. The proposed approach is applied to study of dynamic connectivity patterns of sea surface temperatures in the Pacific Ocean and reveals meaningful edges.
@@ 1 Introduction
The main tool for inferring conditional dependencies among random functions is functional graphical models (FGMs), where each node represents an entire function instead of a scalar variable. Like for traditional (non-functional) graphical models, there are primarily two approaches for graph estimation with FGMs. The first approach is neighborhood selection, which regresses each variable on all other variables and identifies conditional dependencies based on nonzero regression coefficients. Zapata et al., (2022) and Zhao et al., (2024) adapt this approach for functional data using function-on-function regressions, i.e., linear regression models with a functional response and functional covariates. Related, Wu et al., (2022) estimate multiple graphs simultaneously and add a Frobenius norm penalty to enforce graph similarity. Nonparametric variations that relax linearity and distributional assumptions (see below) are also available (Li and Solea,, 2018; Solea and Dette,, 2022; Solea and Li,, 2022; Lee et al., 2023b, ) and may incorporate external covariates (Lee et al., 2023a, ). However, these methods are exclusively non-Bayesian, do not provide uncertainty quantification, and typically require the selection of multiple tuning parameters.
@@
In general, these methods only estimate a single, time-invariant graph for the multivariate functional data. However, when the data are functions of time, it is often of interest to estimate a dynamic graph, so that the conditional dependencies may change through time. For instance, connectivity patterns among SSTs may differ depending on the phase of the seasonal cycle, motivating dynamic graph estimation for multivariate functional data. This generalization requires careful definition of an appropriate graph. Qiao et al., (2020) propose dynamic graph estimation by (i) smoothing the functional data and (ii) separately estimating sparse graphs at each time $t$. Nodes in this dynamic graph at time $t$ represent functions evaluated only at time $t$. However, the graph at time $t$ is influenced by the functional data observed on the whole time domain as a result of the smoothing in (i), which confuses the interpretation of the dynamic graph and may introduce sensitivity to data far from $t$. Further, this approach requires selection of multiple tuning parameters, including a sparsity parameter for every time $t$, and does not provide uncertainty quantification. Alternatively, Zhang et al., (2021) introduce a Bayesian model that defines an edge when two functions are conditionally dependent at time $t$ given all other functions at all other times. Clearly, this graph definition deviates from traditional and functional graphical models: the conditional dependence here does not correspond exactly to the nodes in the graph. Further, Zhang et al., (2021) require local, basis coefficient-specific shrinkage priors, which can induce unnecessary variability unless additional structure is imposed (Gao and Kowal,, 2024).
@@ 2.1 Bayesian functional graphical models
The goal of functional graphical models (FGMs) is to learn and express the conditional dependencies between random functions. Unlike traditional graphical models, where the nodes usually represent scalar random variables, the FGM nodes represent random functions on their full domain. $\bm{Y}^{i}=\{Y^{i}_{j}(t)\}_{j=1}^{p}$ denote iid replicates for $i=1,\ldots,n$ of $p$ random functions $\{y_{j}\}_{j=1}^{p}$ on the functional domain $t\in\mathcal{T}\subseteq\mathbb{R}$. Since these functions may be observed subject to measurement error, we similarly define $\{X_{j}^{i}(t)\}_{j=1}^{p}$ to be iid realizations of the error-free functions $\{x_{j}\}_{j=1}^{p}$. The conditional dependencies among the $p$ random functions are represented by an undirected graph $\mathcal{G}(V,E)$ with vertices $\mathcal{V}=\{1,\ldots,p\}$ and undirected edges $\mathcal{E}\subset\mathcal{V}\times\mathcal{V}$. An edge $(j_{1},j_{2})$ exist if and only if the corresponding two functions are not conditionally independent. Specifically, we define conditional independence for functional data as in Zhu et al., (2016): letting $A,B,C\subset\{1,\ldots,p\}$ we say that $\bm{x}_{A}$ and $\bm{x}_{B}$ are conditionally independent given $\bm{x}_{C}$ if and only if the conditional distribution $[\bm{x}_{A}\mid\bm{x}_{B},\bm{x}_{C}]$ equals the conditional distribution $[\bm{x}_{A}\mid\bm{x}_{C}]$.
@@
The fundamental idea of Gaussian graphical models is that conditional independence corresponds to sparsity in the precision matrix. Let $\bm{\Omega}_{j_{1},j_{2}}=\{\omega_{j_{1}k_{1},j_{2}k_{2}}\}_{k_{1},k_{2}=1}^{K}$ denote the $(j_{1},j_{2})$th block of $\bm{\Omega}$ for $j_{1},j_{2}=1,\ldots,p$, so each block is $K\times K$. Under the Gaussianity assumption (3), it is easy to show that a zero block $\bm{\Omega}_{j_{1},j_{2}}=\bm{0}_{K\times K}$ is equivalent to the conditional independence between $\bm{c}_{j_{1}}$ and $\bm{c}_{j_{2}}$ in (4). Therefore, we deduce that, based on (4), a zero block $\bm{\Omega}_{j_{1},j_{2}}=\bm{0}_{K\times K}$ is equivalent to the absence of an edge between nodes $j_{1}$ and $j_{2}$ in the functional space $\mathcal{G}(V,E)$. In other words, learning the edge set $E$ of the graph $\mathcal{G}(V,E)$ is equivalent to learning the nonzero blocks of $\bm{\Omega}$. In order to recover edges of the graph $\mathcal{G}$ in the functional space, we must learn the block-sparsity structure of $\bm{\Omega}$.
@@
Crucially, this prior incorporates a hierarchical block-wise sparsity structure for $\bm{G}$. As $\pi_{j_{1}j_{2}}$ tends toward zero, the edges in the $(j_{1},j_{2})$th block of $\bm{G}$ are increasingly likely to be all zeros. This corresponds to a lack of an edge $(j_{1},j_{2})$ in the functional space graph $\mathcal{G}$. Unlike previous approaches that use a single weight for the spike-and-slab mixture components across all edges (Wang,, 2015), our mixture weights $\pi_{j_{1}j_{2}}$ are (i) block-specific and (ii) assigned hierarchical Beta priors, thus learned from the data. This prior (6)–(7) is carefully designed for data-adaptive, block-wise sparsity in $\bm{G}$, which leads to sparse and accurate estimation of the function space graph $\mathcal{G}$ (see Section 3).
Section 2.2.3 provides supplementary analysis of the prior based on Monte Carlo samples. The hyperparameters $\alpha$ and $\beta$ in the Beta distribution control the level of sparsity in the graph $\bm{G}$, which is then used to infer the graph $\mathcal{G}$. When the mean is higher, the resulting graphs in the basis space (Figure 1) and the functional space (Figure 8) are denser. In the simulation study and the real data analysis, we assume that the on-diagonal blocks are always fully connected and fix $\pi_{0}=1$ for simplicity; results are not sensitive to this choice. Lastly, we specify an inverse-Gamma prior for $\sigma_{\epsilon}^{2}$ with shape $\alpha_{\sigma}$ and rate $\beta_{\sigma}$.
@@ 2.2.2 Posterior inference
An alternative approach for functional graph estimation is to convert each posterior sample of $\bm{G}^{(s)}$ into a function space graph $\mathcal{G}^{(s)}$, and then compute the posterior median of $\mathcal{G}^{(s)}$ directly. However, the posterior edge inclusion probabilities under this approach are typically much larger than for the previous approach, and often much larger than 0.5. Thus, higher threshold probabilities are required to obtain sparse functional graph estimates. We prefer the previous approach because it more directly leverages the hierarchical block-wise sparsity prior for $\bm{G}$ and produces sparse yet accurate graph estimates for $\mathcal{G}$.
@@ 2.2.3 Choice of hyperparameters
Bayesian graphical models commonly set the prior marginal edge inclusion probability to $2/(p-1)$, which implies that the prior expected number of edges is $p$. Accordingly, we set the prior belief of edge inclusion probabilities of graph $\bm{G}^{(s)}$ in the coefficient space to $2/(p-1)$, corresponding to an expected number of edges equal to $pK^{2}$. As in Wang, (2015) and Peterson et al., (2020), the prior distribution of graph $\bm{G}$ is not analytically available because of intractable normalizing constants. A complete form of the full prior with the normalizing constants is included in Appendix A. In order to investigate whether the normalizing constants in our new model dominate the prior and drive the prior to the extreme, we evaluate the prior edge inclusion probability based on Monte Carlo sampling from the prior.
Figure 1: Prior edge inclusion probability (y-axis) versus mean of the Beta prior, $\alpha/(\alpha+\beta)$, (x-axis) for $p=5$ (top left), $p=10$ (top right), $p=15$ (bottom left), and $p=20$ (bottom right). The dotted gray line is when the two axes are equal.
First, we evaluate the influence of $\alpha$ and $\beta$ on the marginal prior edge inclusion probability. We fix $\lambda=1$, $\nu_{0}=0.02$ and $h=50$ ($v_{1}=hv_{0}$) as commonly recommended (Wang,, 2015; Peterson et al.,, 2020; Osborne et al.,, 2022). We also fix $\alpha=2$ and $K=5$. The value of $\beta$ varies so that the prior mean of the beta distribution kernel for each $\pi_{j_{1},j_{2}}$ in (7) takes values in $\{1/(p-1),2/(p-1),3/(p-1),4/(p-1),5/(p-1)\}$. The study is carried out on $p=(5,10,15,20)$. Figure 1 shows the prior edge inclusion probabilities estimated via Monte Carlo sampling. We conclude that: 1) the prior edge inclusion probability is not dominated by the normalizing constant; 2) similar to Wang, (2015), the prior edge inclusion probability is lower than the prior mean of $\pi_{j_{1},j_{2}}$; and 3) the influence of normalizing constants on the prior edge inclusion probability is greater when $p$ is smaller, but the differences among different values of $p$ are negligible. Figure 1 can be used as a guide to choose parameters that reflect the preferred edge inclusion probability. In the example of $p=15$ the prior edge inclusion probability is slightly below $2/(p-1)$ when the prior mean of $\pi_{j_{1},j_{2}}$ is $3/(p-1)$. Therefore, when $p=15$, $\alpha=2$ and $\beta=7$ is a reasonable choice.
@@ 3.3 Parameter settings
We use $K=5$ B-splines of order 3 to fit the model, which provides accurate estimation of the underlying true smooth curves. Note that this basis differs from the polynomial basis used to generate the data. For the hyperparameters in the joint prior (5), we follow the guidelines in Section 2.2.3 and set $\lambda=1$, $v_{0}=0.02$, $h=50$ ($v_{1}=hv_{0}$), $\alpha=2$ and $\beta=7$. To initialize the MCMC, the precision matrix and graphs are all set to the identity matrix and the block-wise edge inclusion probabilities are set to 0.5. Our simulation study shows that the model is robust to the initialization of the basis coefficients and the changepoint $\tau$. Results reported below were obtained by running MCMC chains for 5000 total iterations with a burn-in of 3000 iterations. The models took on average about 4 minutes on a Mac computer with a 16 GB Apple M1 Pro chip. MCMC convergence was checked by the Geweke diagnostic test (Geweke et al.,, 1991) on the key parameters $\{\bm{c}^{i(s)},\sigma_{\epsilon}^{(s)},\tau\}$.
@@ 3.5 Learning the block-sparsity
As suggested by our simulation results, our proposed model, together with the choice of hyper-parameters, leads to sparse posterior estimation of the $p$-by-$p$ graphs in the functional space. Our prior on $\bm{G}$ does not require specific edges to be included or excluded, but rather gives a common marginal (over all other parameters) probability of inclusion to all edges. These parameters induce group shrinkage and enable information sharing among edges within a block. If the data indicate that most of the edges in a block are excluded, then we expect the posterior of $\pi_{j_{1}j_{2}}$ to concentrate on small values, which will further discourage the inclusion of other edges in that block.
@@
$p(g_{j_{1}k_{1},j_{2}k_{2}}=1\mid\bm{\pi},\bm{\Omega})=\dfrac{N(\omega_{j_{1}k_{1},j_{2}k_{2}}\mid 0,v_{1}^{2})\pi_{j_{1},j_{2}}}{N(\omega_{j_{1}k_{1},j_{2}k_{2}}\mid 0,v_{1}^{2})\pi_{j_{1}j_{2}}+N(\omega_{j_{1}k_{1},j_{2}k_{2}}\mid 0,v_{0}^{2})(1-\pi_{j_{1}j_{2}})},$
which is a combination of $\pi_{j_{1},j_{2}}$ and the likelihood. In our simulation study, the posterior mean of $\pi_{j_{1},j_{2}}$ was about 0.08 for the zero blocks and 0.19 for the nonzero blocks. Edges with a high likelihood in the slab component have a high posterior marginal inclusion probability, which results in an estimated edge in the functional space. To illustrate this point, we calculated the Frobenuius norm of estimated precision matrix ^$\bm{\hat{}}{\bm{\Omega}}$ in each block, and compared the distribution of these Frobenius norms between zero and nonzero blocks (see Appendix B). There is a clear separation between the two groups. The density from the zero blocks is highly concentrated around zero, especially compared to the density from the nonzero blocks.